# Geometry-Aware Diffusion Models for Multiview Scene Inpainting

### v1

@@ Abstract
In this paper, we focus on 3D scene inpainting, where parts of an input image set, captured from different viewpoints, are masked out. The main challenge lies in generating plausible image completions that are geometrically consistent across views. Most recent work addresses this challenge by combining generative models with a 3D radiance field to fuse information across viewpoints. However, a major drawback of these methods is that they often produce blurry images due to the fusion of inconsistent cross-view images. To avoid blurry inpaintings, we eschew the use of an explicit or implicit radiance field altogether and instead fuse cross-view information in a learned space. In particular, we introduce a geometry-aware conditional generative model, capable of inpainting multi-view consistent images based on both geometric and appearance cues from reference images. A key advantage of our approach over existing methods is its unique ability to inpaint masked scenes with a limited number of views (i.e., few-view inpainting), whereas previous methods require relatively large image sets for their 3D model fitting step. Empirically, we evaluate and compare our scene-centric inpainting method on two datasets, SPIn-NeRF and NeRFiller, which contain images captured at narrow and wide baselines, respectively, and achieve state-of-the-art 3D inpainting performance on both. Additionally, we demonstrate the efficacy of our approach in the few-view setting compared to prior methods. Our project page is available at https://geomvi.github.io.
@@ 1 Introduction
Image inpainting is a long-standing problem in computer vision and graphics [3]. Unlike unconditional generation, inpainting is constrained by the conditioning input, requiring the output to be visually plausible and to match the partial content. The problem of inpainting 3D scenes has recently grown in popularity (e.g., [42, 56]) due to the advent of powerful novel view synthesis (NVS) models. Such models are usually implemented as scene models capable of differentiable rendering, e.g., neural radiance fields (NeRFs) [40] or Gaussian splatting [25]. For both NVS and 3D inpainting, the input scene is implicitly represented through a posed set of images; hence, we may consider the equivalent problem of “multiview inpainting”, which provides a natural interface between 2D image inpainting and 3D NVS.
@@
A common approach, often used in NeRF editing (e.g., [72]), is to fuse information across views, via 3D scene models designed for NVS. A cyclic process, such as iterative dataset update [18], is then used, whereby images are edited, used to fit the NVS model, and NVS renders (which combine information across views) assist with new edits. While this approach ensures consistency via the integrated 3D radiance field, it has a few shortcomings: (i) a tendency towards blurry outputs and (ii) a reliance on the NVS scene model, which presents inherent difficulties of its own. The blur in (i) is, in part, due to the fusion happening in pixel space, while the downsides of (ii) include the need for accurate cameras, with sufficient view coverage.
In this work, we resolve these problems by fusing cross-view information in a learned space, during a generative diffusion process, and eschewing the necessity of an NVS model, though we can optionally choose to use one as a separate post-fitting step. We further reduce the tendency of generative inpainters to over-hallucinate, a common cause of 3D inconsistencies, by conditioning them on the 3D scene structure. In particular, we devise a geometry-aware conditional diffusion model, capable of inpainting multiview-consistent images based on geometric and appearance cues from reference images. One key capability of our model is the handling of uncertain or partial scene information (e.g., missing due to occlusions and viewpoint changes). We integrate our model into a 3D scene inpainting algorithm, performing 3D-consistent propagation of image content across views. Unlike NeRF-based inpainters, which fuse inconsistencies in appearance space and thus induce blur, our method fuses cross-view information via the generative model, resulting in sharper outputs even when a NeRF is fit to our final inpainted results. In this paper, we target three main inpainting tasks: narrow-baseline object removal, wide-baseline scene completion, and few-view inpainting. Most previous work has been focused on the first task, the second one is a recent addition, and the third task has not been extensively explored with recent innovations. Fig. 1 provides a visualization of each of our target tasks. We evaluate our multiview inpainter on two datasets, SPIn-NeRF [42] and NeRFiller [72], which contain narrow and wide baselines, respectively, and achieve state-of-the-art 3D inpainting performance on both. Please see our project page for visualizations of results: https://geomvi.github.io.
@@ 2 Related Work
Many 3D inpainters focus primarily on object removal [70, 36, 42, 73, 65]. In contrast, our method can insert additional content and perform scene completion. RenderDiffusion [1] is capable of 3D-aware inpainting, but focuses on weak supervision (utilizing only 2D supervision) and simpler scenes. SIGNeRF [13] focuses on localized translation and object insertion, rather than general 3D inpainting. Chen et al. [9] focuses on the role of the inpainting mask for deterministic object removal, but only handles forward-facing scenes. Gaussian Grouping [77] builds segmentation features directly into the radiance field, improving the ease of a variety of editing tasks. For general inpainting, most methods build on 3D radiance fields [8, 44, 48, 34, 30]. Several methods (e.g., [44, 48, 8]) apply score distillation sampling (SDS) [47, 66] to utilize 2D diffusion priors. In contrast, our method inpaints the image set directly. Similar to our work, several recent inpainters also rely on altering diffusion models, specializing them for multiview inpainting [5, 44]. For our evaluation, we utilize the scene-centric settings of NeRFiller [72] and SPIn-NeRF [42], which covers a variety of scene types and camera baselines.
@@
Reference-Conditioned Generative Editing. The advent of diffusion foundation models for 2D image generation is promising for 3D inpainting. However, independent inpaintings of multiview images are inconsistent, particularly due to the high diversity of stochastic inpainters. To faciliate editing, a variety of methods have been devised to encourage consistency across generations (e.g., [63, 89, 2, 54]), usually by sharing features across diffusion processes. However, these methods are only able to enforce semantic consistency, which is insufficient to ensure the precise (i.e., pixel-level) coherence required for generating 3D-consistent content. Separately, conditional image generators can perform novel view synthesis (NVS) by mapping observed reference images and target camera parameters to a new view (e.g., [33, 57, 75, 16, 79, 64, 78, 6, 80]) by training on multiview data. However, these methods are meant for generative NVS, rather than inpainting existing 3D scenes with coherent content. While our conditional diffusion model is reminiscent of this, addressing the entirety of the NVS problem is not necessary for 3D inpainting; instead, we assemble reference information (from multiple source views) in the image coordinate frame of the target view (i.e., no camera information is sent to the model). Hence, our inpainter does not need to learn how to perform complex 3D transforms of image content; instead, it needs only to learn to utilize the information projected from the source views, entirely in 2D image space. Finally, there is the notion of “reference-based inpainting” [85, 83, 88], where one image is inpainted via content from a reference image of the scene (though they may also be required to handle transforms beyond viewpoint, such as lighting changes). While our approach does use reference information to condition the inpainting, our method is more similar to other NeRF-based methods, in that we operate on full multiview image sets, where all views are initially masked, rather than just a single reference-target pair.
@@
Iterative 3D Generation. 3D content generation is typically iterative or autoregressive. Iterative methods, common in text-to-3D [47, 38] and instruction-based 3D editing [18], have also been used for 3D inpainting via SDS [47] (e.g., [48, 44, 8]) and Iterative Dataset Update (IDU) [18] (e.g., [72]). In contrast, autoregressive approaches have been explored in 3D scene generation (e.g., [12, 21]) and generative NVS (e.g., [78, 32, 52, 51, 64]). We introduce an autoregressive approach for inpainting large-baseline scenes.
@@
Figure 2: Overview of our geometric-aware 3D scene inpainter. (a) A visualization of the reference-based geometric cues. Backfaces are always covered by shadow volumes. $T_{i}(I_{i}),F_{i},B_{i},C_{i},\widehat{D}_{i}$ denote the rendered photometric content, front-face mask, back-face mask, shadow mask, and disparity, respectively, for the reference view $i$. (b) A step-by-step visualization of our autoregressive inpainting process. Note that the scene geometry consists of separate meshes for each image, not a single harmonized mesh, as shown here for simplicity. Here, we are only showing one diffusion step for the geometry-aware inpainting model. $\mathcal{E},z_{t},I,M,A_{\mathcal{R}_{i}},\mathcal{G}_{\mathcal{R}_{i}}$ denote the VAE encoder, the diffusion latent at timestep $t$, the masked image latent, the mask, the appearance cues for reference view $i$, and the geometric cues for reference view $i$, respectively.
@@ 3 Method
Setting. We assume as given (i) a set of $N$ views, $I_{i}\in^{H\times W\times 3}$, representing a static scene, and (ii) their corresponding inpainting masks, $M_{i}\in\{0,1\}^{H\times W}$, demarcating the 3D region to be inpainted. While other methods may require additional inputs, such as camera parameters or depths per image, these are optional for our approach (though we can nonetheless use them). Our objective is to jointly inpaint the given views, thus inpainting the 3D scene, ideally in a consistent manner across views.
Overview. Prior works (e.g., [42, 41, 72]) often have two disparate components: (i) a (neural) 3D radiance field, which fuses information across views to ensure consistency, and (ii) a 2D image inpainter, which alters the source views, potentially conditioned on the state of (i) (e.g., [18, 43, 72]). However, this separation has shortcomings: first, (i) fuses information in pixel space (due to the supervision mechanism), leading to blurry results, and second, (ii) is only aware of the other views through (i), meaning any mechanism for enforcing consistency is indirect. In contrast, our algorithm relaxes the requirement for maintaining a complete explicit scene, fuses information in the learned space of a generative model (avoiding the blur of pixel space), and enables direct transfer of appearance information across views, using estimated scene geometry. To attain this, our approach consists of two models: a scene geometry estimator and a geometry-aware image inpainting model. For the former, we use the performant DUSt3R model [68], which efficiently provides dense depth, with or without the presence of cameras, utilizing views (inpainted or not) directly. For the latter, we fine-tune a 2D diffusion-based inpainter, to condition on the other views. However, naively conditioning the inpainter forces it to not only learn inpainting, but also to learn generative NVS (itself a non-trivial task; e.g., [79, 16, 62]), as the model must discern how to transform content from one view to another, using some internal model of the scene geometry. Instead, we use our scene geometry estimator to feed appearance information from other views to our inpainter, by directly projecting information from source views, as well as passing explicit geometric information pertinent to the inpainting (e.g., occlusion).
@@
For the remainder of this section, we provide details on our approach. We first briefly review our scene geometry estimator, based on DUSt3R [68]3.1). Next, we describe our geometry-aware inpainting diffusion model in §3.2, including the reference-based geometric cues guiding each view’s inpainting. Finally, we present the autoregressive procedure used to iteratively inpaint the scene (§3.3). §B of the supplement provides additional methodological details.
@@ 3.1 Scene Geometry Estimation
We utilize DUSt3R [68] for scene geometry reconstruction. DUSt3R can efficiently estimate scene depths and camera parameters from multiview image sets. As we inpaint the scene, we iteratively reapply it to update the scene geometry throughout the process. See §B.1 for details.
@@ 3.2 Geometry-aware Inpainting Diffusion Model
Our goal is to devise a model that performs image inpainting, conditioned on the presence of a multiview image set and scene geometry. The view set may consist of a mixture of masked (non-inpainted) and complete (inpainted) images, which we wish to use to fill in the target image. Thus, this setup is a form of reference-based inpainting, with the additional constraints of a 3D scene structure.
Notation. Formally, let $I$ be the target image to be inpainted, with mask, $M$, and camera, $\Pi$ (i.e., handling a single image in the scene at a time). Each element of the reference view-set, =$\mathcal{R}=\{(I_{i},M_{i},b_{i},\Pi_{i},D_{i})\}_{i}$, consists of an image ($I_{i}$), mask ($M_{i}$), indicator of whether $I_{i}$ has been inpainted or filled in ($b_{i}$), camera parameters ($\Pi_{i}$), and depth map ($D_{i}$). When image $I_{i}$ has already been inpainted (so $b_{i}=1$), we simply set $M_{i}$ to nullity (i.e., all parts of $I_{i}$ are trustworthy). Camera and depth information, in our case, come from our geometry estimator (§3.1). Thus, based on the current state of the scene, our inpainter, $f$, obtains $\widehat{I}=f(I,M|\mathcal{R})$, which can later be used in the reference view-set of future inpaintings (see §3.3).
@@
The core of our conditioning approach is to take advantage of geometric knowledge of the scene. In particular, using scene geometry, we project appearance and geometric information from references into the target camera’s image plane; this alleviates the need for the network to learn geometry estimation and view transforms (as in NVS), saving its capacity for inpainting. Given depth and cameras, from §3.1, this is a simple projective mapping. However, three issues remain: (i) selecting cues to pass to $f$, (ii) fusing cues across views, and (iii) training the inpainter to handle errors in estimated scene geometry. We first consider (i), then discuss (ii) in §3.2.2 and (iii) in §3.2.3.
@@
Preliminaries. For each reference image, $I_{i}$, let $T_{i}$ be the projective transform from the $i$th camera to the frame of the target frame ($I$). This is performed by constructing a mesh, $S_{i}$, from $D_{i}$, assigning cues as nodal attributes, and rendering $S_{i}$ into the target frame via $\Pi$ (see §B.2 for details).
@@
Appearance Cues. We provide $f$ with two appearance cues: (a) direct reference pixels and (b) a stylistic hint. For (a), we pass $T_{i}(I_{i})$ instead of $I_{i}$, giving the inpainter direct access to view-aligned pixel colours for localized convolutional processing. However, with large camera baselines, projected information from distant viewpoints may be insufficient (e.g., only the back-face triangles of an object are visible). To ensure a stylistic harmony, we introduce (b), an optional “hint image”, $H$, which is not projected through the scene geometry but rather provides global appearance characteristics. However, we still want to inform $f$ of its appearance, to ensure a stylistically harmonized inpainting. When used, we set $H$ as the furthest inpainted image from $I$ in $\mathcal{R}$. See §B.4 for details. Mathematically, let $A_{\mathcal{R}}=\{T_{i}(I_{i})\}_{i}\cup\{H\}$ be the set of appearance cues.
@@
Geometric Cues. We next obtain a set of geometry-based cues, which control the trustworthiness of photometric reference content, illustrated in Fig. 2 (a). Specifically, per reference, we compute a (i) front-face mask, (ii) back-face mask, (iii) normalized inverse depth (ordering), and (iv) shadow mask, all in the target coordinate frame. The mesh $S_{i}$ has front and back faces, indicating the validity of photometric content (i.e., the former has valid appearance from $I_{i}$, but the latter merely implicates the presence of geometry). Thus, (i), (ii), and (iii) are rendered from $S_{i}$, denoted $F_{i}$, $B_{i}$, and $\widehat{D}_{i}=T_{i}(D_{i})$, respectively. For (iv), $S_{i}$ implies a “shadow volume” [39, 17] with respect to $\Pi_{i}$, representing the 3D space hidden from the reference camera (see Fig. 2 and §B.3 for details), which is rendered into the target view, written $C_{i}$. From the target view, any pixel with a camera ray hitting the shadow volume is explicitly uncertain – there could be content there or not. Overall, these maps are assembled into our geometric cue set, $\mathcal{G}_{\mathcal{R}}=\{F_{i},B_{i},\widehat{D}_{i},C_{i}\}_{i}$, which will be fed into our diffusion model. Importantly, these cues also enable a hierarchical, confidence-based fusion of reference information, based on the uncertainty induced by the geometric structure.

### v2

@@ Abstract
In this paper, we focus on 3D scene inpainting, where parts of an input image set, captured from different viewpoints, are masked out. The main challenge lies in generating plausible image completions that are geometrically consistent across views. Most recent work addresses this challenge by combining generative models with a 3D radiance field to fuse information across a relatively dense set of viewpoints. However, a major drawback of these methods is that they often produce blurry images due to the fusion of inconsistent cross-view images. To avoid blurry inpaintings, we eschew the use of an explicit or implicit radiance field altogether and instead fuse cross-view information in a learned space. In particular, we introduce a geometry-aware conditional generative model, capable of multi-view consistent inpainting using reference-based geometric and appearance cues. A key advantage of our approach over existing methods is its unique ability to inpaint masked scenes with a limited number of views (i.e., few-view inpainting), whereas previous methods require relatively large image sets for their 3D model fitting step. Empirically, we evaluate and compare our scene-centric inpainting method on two datasets, SPIn-NeRF and NeRFiller, which contain images captured at narrow and wide baselines, respectively, and achieve state-of-the-art 3D inpainting performance on both. Additionally, we demonstrate the efficacy of our approach in the few-view setting compared to prior methods. Our project page is available at https://geomvi.github.io.
@@ 1 Introduction
Image inpainting is a long-standing problem in computer vision and graphics [3]. Unlike unconditional generation, inpainting is constrained by the conditioning input, requiring a visually-plausible output that matches the partial content. The problem of inpainting 3D scenes has recently grown in popularity (e.g., [43, 59]) due to the advent of powerful novel view synthesis (NVS) models. Such models are usually implemented as scene models capable of differentiable rendering, e.g., neural radiance fields (NeRFs) [41] or 3D Gaussian splatting (3DGS) [26]. For both NVS and 3D inpainting, the input scene is implicitly represented through a posed set of images; hence, we may consider the equivalent problem of “multiview inpainting”, which provides a natural interface between 2D image inpainting and 3D NVS.
@@
A common approach, often used in NeRF editing (e.g., [75]), is to fuse information across views, via 3D radiance fields. A cyclic process, such as iterative dataset update (IDU) [18], is then used, whereby images are edited, used to fit the NVS model, and NVS renders (which combine information across views) assist with new edits. While this approach ensures consistency via the 3D radiance field, it has a few shortcomings: (i) a tendency towards blurriness, partly due to the fusion happening in pixel space, and (ii) reliance on the radiance field, which requires accurate camera parameters with sufficient view coverage.
In this work, we resolve these problems by fusing cross-view information in a learned space, during a generative diffusion process, and eschewing the necessity of a 3D radiance field, though we can optionally choose to use one as a separate post-fitting step. We further reduce the tendency of generative inpainters to over-hallucinate, a common cause of 3D inconsistencies, by conditioning them on the 3D scene structure. In particular, we devise a geometry-aware conditional diffusion model, capable of inpainting multiview-consistent images based on geometric and appearance cues from reference images. One key capability of our model is the handling of uncertain or partial scene information (e.g., missing due to occlusions and viewpoint changes). We integrate our model into a 3D scene inpainting algorithm, performing 3D-consistent propagation of image content across views. Unlike NeRF-based inpainters, which fuse inconsistencies in appearance space and thus induce blur, our method fuses cross-view information via the generative model, resulting in sharper outputs even when a NeRF is fit to our final inpainted results. In this paper, we target three main inpainting tasks: narrow-baseline object removal, wide-baseline scene completion, and few-view inpainting. Most previous work has been focused on the first task, the second one is a recent addition, and the third task has not been extensively explored with recent innovations. Fig. 1 provides a visualization of each of our target tasks. We evaluate our multiview inpainter on two datasets, SPIn-NeRF [43] and NeRFiller [75], which contain narrow and wide baselines, respectively, and achieve state-of-the-art 3D inpainting performance on both. Please see our project page for visualizations of results: https://geomvi.github.io.
@@ 2 Related Work
2D Priors for 3D Inpainting. 2D inpainting has been widely explored [50]. Current state-of-the-art methods leverage conditional diffusion models [55, 85, 24], which exhibit high-quality and diverse generations. However, direct application of independent 2D inpainters for 3D inpainting leads to 3D inconsistencies [43, 42]. Yet, despite the challenges in multiview inconsistency, recent methods [75, 49, 45, 8, 31] have explored combining diffusion-based 2D inpainters with explicit or implicit 3D radiance fields to exploit their powerful generative prior. To preserve this prior while enforcing greater 3D consistency, we further constrain a diffusion-based inpainter using scene geometry.
@@
Many 3D inpainters focus primarily on object removal [73, 37, 43, 76, 68]. In contrast, our method can insert additional content and perform scene completion. RenderDiffusion [1] enables 3D-aware inpainting, but relies on weak supervision (utilizing only 2D supervision) and is limited to simpler scenes. SIGNeRF [13] specializes in localized translation and object insertion, rather than general 3D inpainting. Chen et al. [9] examines the impact of the inpainting mask for deterministic object removal, but is restricted to forward-facing scenes. Gaussian Grouping [80] integrates segmentation features directly into the radiance field, facilitating various editing tasks. For general inpainting, most methods build on 3D radiance fields [8, 45, 49, 35, 31], with several approaches (e.g., [45, 49, 8]) leveraging score distillation sampling (SDS) [48, 69] to incorporate 2D diffusion priors. However, these methods typically require a relatively dense view coverage for 3D model fitting. In contrast, our method inpaints the image set directly, allowing it to operate effectively even with sparse view coverage. Similar to our work, several recent inpainters also alter diffusion models, specializing them for multiview inpainting [5, 45]. For our evaluation, we utilize the scene-centric settings of NeRFiller [75] and SPIn-NeRF [43], which covers a variety of scene types and camera baselines.
@@
Reference-Conditioned Generative Editing. A variety of methods have been devised to encourage consistency across multiple generations (e.g., [66, 92, 2, 57]), usually by sharing features across diffusion processes. However, these methods only enforce semantic consistency, which is insufficient for precise, pixel-level coherence required for 3D-consistent content. Separately, conditional image generators can perform NVS by mapping observed reference images and target camera parameters to a new view (e.g., [34, 60, 78, 16, 82, 67, 81, 6, 83]) by training on multiview data. However, these methods are meant for generative NVS, rather than inpainting existing 3D scenes coherently. While our conditional diffusion model is reminiscent of this, addressing the entirety of the NVS problem is not necessary for 3D inpainting; instead, we assemble reference information (from multiple source views) in the image coordinate frame of the target view (i.e., no camera information is sent to the model). Hence, our inpainter avoids learning complex 3D transforms (e.g., triangulating and projecting between frames). Instead, it learns to utilize projected reference information entirely in 2D, while accounting for simple 3D constraints, like relative depth ordering of generated content with respect to existing content. Finally, “reference-based inpainting” [88, 87, 91] uses one reference image to inpaint another, sometimes handling transforms beyond viewpoint, like lighting changes. While our method also leverages reference information, it aligns more with NeRF-based approaches, operating on full multiview image sets where all views are initially masked, rather than a single reference-target pair.
@@
Iterative 3D Generation. 3D content generation is typically iterative or autoregressive. Iterative methods, common in text-to-3D [48, 39] and instruction-based 3D editing [18], have also been used for 3D inpainting via SDS [48] (e.g., [49, 45, 8]) and IDU [18] (e.g., [75]). In contrast, autoregressive approaches have been explored in 3D scene generation (e.g., [12, 21]) and generative NVS (e.g., [81, 33, 54, 53, 67]). We introduce an autoregressive approach for inpainting large-baseline scenes.
@@
Figure 2: Overview of our geometric-aware 3D scene inpainter. (a) A visualization of the reference-based geometric cues. Back-faces are always covered by shadow volumes. $T_{i}(I_{i}),F_{i},B_{i},C_{i},\widehat{D}_{i}$ denote the rendered photometric content, front-face mask, back-face mask, shadow mask, and disparity, respectively, for the reference view $i$. (b) A step-by-step visualization of our autoregressive inpainting process. Note that the scene geometry consists of separate meshes for each image, not a single harmonized mesh, as shown here for simplicity. Here, we are only showing one diffusion step for the geometry-aware inpainting model. $\mathcal{E},z_{t},I,M,A_{\mathcal{R}_{i}},\mathcal{G}_{\mathcal{R}_{i}}$ denote the VAE encoder (which maps an image to the latent space of Stable Diffusion), the diffusion latent at timestep $t$, the masked image latent, the mask, the appearance cues for reference view $i$, and the geometric cues for reference view $i$, respectively; see §F.3 for an example of autoregressive steps.
@@ 3 Method
Setting. As with prior work [43], we assume as given a set of $N$ views, $I_{i}\in^{H\times W\times 3}$, representing a static scene, and their corresponding inpainting masks, $M_{i}\in\{0,1\}^{H\times W}$, demarcating the 3D region to be inpainted. While other methods may require additional inputs, such as camera parameters or depths per image, these are optional for our approach (though we can nonetheless use them). Our objective is to jointly inpaint the given views, thus inpainting the 3D scene, ideally in a consistent manner across views.
Overview. Prior works (e.g., [43, 42, 75, 49, 45, 8]) often have two disparate components: (i) an implicit or explicit 3D radiance field (e.g., NeRF [41] and 3DGS [26]), which fuses information across views to ensure consistency, and (ii) a 2D image inpainter, which alters the source views, potentially conditioned on the state of (i) (e.g., [18, 44, 75]). However, this separation has shortcomings: first, (i) fuses information in pixel space (due to the supervision mechanism), leading to blurry results, and second, (ii) is only aware of the other views through (i), meaning any mechanism for enforcing consistency is indirect. In contrast, our algorithm relaxes the need for a radiance field, fuses information in the learned space of a generative model (avoiding the blur of pixel space), and enables direct appearance transfer across views, using estimated scene geometry. To attain this, our approach consists of two models: a scene geometry estimator and a geometry-aware inpainting model. For the former, we use the performant DUSt3R [71], which efficiently provides dense depth, with or without the presence of camera poses, utilizing views (inpainted or not) directly. For the latter, we fine-tune a latent 2D diffusion-based inpainter, to condition on the other views. However, naively conditioning the inpainter forces it to learn both inpainting and generative NVS (itself a non-trivial task [82, 16, 65]), by learning to map one view to another using an internal scene model. Instead, we use our scene geometry estimator to feed appearance information from other views to our inpainter, by directly projecting information from source views, as well as passing explicit geometric information pertinent to the inpainting (e.g., occlusion).
@@
For the remainder of this section, we provide details on our approach. We first briefly review our scene geometry estimator, based on DUSt3R [71]3.1). Next, we describe our geometry-aware inpainting diffusion model in §3.2, including the reference-based geometric cues guiding each view’s inpainting. Finally, we present the autoregressive procedure used to iteratively inpaint the scene (§3.3). §B provides additional methodological details.
@@ 3.1 Scene Geometry Estimation
We utilize DUSt3R [71] for scene geometry reconstruction. DUSt3R can efficiently estimate scene depths and camera parameters from multiview image sets. As we inpaint the scene, we iteratively reapply it to update the scene geometry throughout the process; see §B.1 for details.
@@ 3.2 Geometry-aware Inpainting Diffusion Model
Our goal is to devise an image inpainting model, conditioned on a multiview image set and scene geometry. The view set may include both masked (non-inpainted) and complete (inpainted) images, used to inpaint the target image. Thus, this setup is a form of reference-based inpainting, with the additional constraints of a 3D scene structure.
Notation. Formally, let $I$ be the target image to be inpainted (i.e., inpainting one image), with mask, $M$, and camera parameters, $\Pi$ (i.e., extrinsics and intrinsics). Each element of the reference view-set, =$\mathcal{R}=\{(I_{i},M_{i},b_{i},\Pi_{i},D_{i})\}_{i}$, consists of an image ($I_{i}$), mask ($M_{i}$), indicator of whether $I_{i}$ has been inpainted ($b_{i}$), camera parameters ($\Pi_{i}$), and depth map ($D_{i}$). When $b_{i}=1$, we simply set $M_{i}$ to nullity (i.e., all parts of $I_{i}$ are trustworthy). Camera and depth information come from our geometry estimator (§3.1). Thus, based on the current state of the scene, our inpainter, $f$, obtains $\widehat{I}=f(I,M|\mathcal{R})$, which can later be used in the reference view-set of future inpaintings (see §3.3).
@@
Our conditioning approach leverages geometric knowledge of the scene. In particular, using scene geometry, we project appearance and geometric information from references into the target camera’s image plane; this alleviates the need for the network to learn geometry estimation and view transforms (as in NVS), saving its capacity for inpainting. Given depth and cameras, from §3.1, this is a simple projective mapping. However, three issues remain: (i) selecting cues, (ii) fusing cues across views, and (iii) training the inpainter to handle errors in estimated scene geometry. We first consider (i), then discuss (ii) in §3.2.2 and (iii) in §3.2.3.
@@
Preliminaries. For each reference image, $I_{i}$, let $T_{i}$ be the projective transform from the $i$th camera to the target frame ($I$). This is performed by constructing a mesh, $S_{i}$, from $D_{i}$, assigning cues as nodal attributes, and rendering $S_{i}$ into the target frame via $\Pi$ (see §B.2 for details).
@@
Appearance Cues. We provide the inpainter, $f$, with two appearance cues: (a) direct reference pixels and (b) a stylistic hint. For (a), we pass $T_{i}(I_{i})$ instead of $I_{i}$, giving the inpainter direct access to view-aligned pixel colours for localized convolutional processing. However, with large camera baselines, projected information from distant viewpoints may be insufficient (e.g., only the back-face triangles of an object are visible). To ensure a stylistic harmony, we introduce (b), an optional “hint image”, $H$, which is not projected through the scene geometry but rather provides global appearance characteristics. When used, we set $H$ as the furthest inpainted image from $I$ in $\mathcal{R}$ (see §B.4). We denote our appearance cues as $A_{\mathcal{R}}=\{T_{i}(I_{i})\}_{i}\cup\{H\}$.
@@
Geometric Cues. We next obtain a set of geometric cues, controlling the reliability of photometric reference content, illustrated in Fig. 2 (a). Specifically, per reference, we compute a (i) front-face mask, (ii) back-face mask, (iii) normalized inverse depth (ordering), and (iv) shadow mask, all in the target coordinate frame. The mesh $S_{i}$ has front- and back-faces, indicating the validity of photometric content (i.e., the former has valid appearance from $I_{i}$, but the latter merely implicates the presence of geometry). Thus, (i), (ii), and (iii) are rendered from $S_{i}$, denoted $F_{i}$, $B_{i}$, and $\widehat{D}_{i}=T_{i}(D_{i})$, respectively. For (iv), $S_{i}$ implies a “shadow volume” [40, 17] with respect to $\Pi_{i}$, representing the 3D space hidden from the reference camera (see Fig. 2 and §B.3 for details), which is rendered into the target view, written $C_{i}$. From the target view, any pixel with a camera ray hitting the shadow volume is explicitly uncertain – there could be content there or not. These maps form our geometric cue set, $\mathcal{G}_{\mathcal{R}}=\{F_{i},B_{i},\widehat{D}_{i},C_{i}\}_{i}$, another input to our diffusion model. Importantly, these cues also enable a hierarchical, confidence-based fusion of reference information, based on the uncertainty induced by the geometric structure.